Suppose f(x) 24x + 14, f( - 1) = 3, and f( - 1) = 5. Then f(x) = Check Answer
Answer #1

Given fW) = 24x+14 four s fleredent ca & (x) = {(240414) de = 24 Judxi Sikde Seth dua khal Isadeaux . nt), 224.22).+ 14x f=128xdeot 14 sade at sodu created the a (24) +5x fim, = 4x²+70+5 atce -12 3 P(-1)=5 PC-1) = 46-19P+ 7(-1) ++56-12 + c 5=-4+7